# A Tutte-type canonical decomposition of 3- and 4-connected graphs

### v2

@@ Abstract.
As an application of the decomposition for 4-connectivity, we obtain a new theorem characterising all quasi-4-connected vertex-transitive finite graphs as essentially quasi-5-connected or on a short explicit list of graphs.
@@
4-connected, decomposition, Tutte, canonical, double wheel, 4-separation, 4-separator, tetra-separation, tetra-separator
@@ 1. Introduction
Background. To solve ˜1.1 for $k=0$ and $k=1$, one simply cuts along all $k$-separators of the graph. But already for $k=2$, this approach fails. Indeed, different 2-separators can cross in that they cut through each other, as happens in a cycle. Then if we cut at one of the 2-separators, we loose the other. Tutte realised that the solution is to cut the graph at all 2-separators that are totally-nested in that they are not crossed by any 2-separators, and indeed no conflicts arise when cutting in this way. The resulting pieces of Tutte’s decomposition are 3-connected, cycles or $K_{2}$’s. All decompositions for $k\leqslant 2$ can be described in terms of tree-decompositions. For instance, Tutte’s selection of 2-separators gives a recipe for a tree-decomposition with all adhesion-sets of size $2$ and all whose torsos [^1] A torso of a tree-decomposition is obtained from a bag by making each of its adhesion-sets into a clique. are 3-connected, cycles or $K_{2}$’s.
@@
There are $k$-connected graphs which just slightly fail to be $(k+1)$-connected, but all whose $k$-separators are crossed. For example, if $G$ is $k$-connected and the neighbourhoods of its vertices are precisely the $k$-separators of $G$, then the $k$-separators around neighbouring vertices cross, but $G$ fails to be $(k+1)$-connected. Such $G$ exist for all $k\geqslant 3$: take a circular saw as in Figure˜1.1, for instance. We believe that such graphs should not have to be decomposed any further at level $k$.
Figure $1.$ A circular saw has vertex-set $\mathbb{Z}_{n}\times\mathbb{Z}_{2}$ with $n\gg k$ and every vertex $(v,0)$ sends edges to the $k$ vertices $(v,1),\ldots,(v+k-1,1)$. It is $k$-connected, and when $k\geqslant 3$ its neighbourhoods of vertices are precisely its $k$-separators (Lemma˜14.2).
@@
We consider the $k$-connected graph $G$ depicted in Figure˜1.2 if $k$ is odd and in Figure˜1.3 if $k$ is even. Every $k$-separator of $G$ is formed from a red set, a non-adjacent blue set, and if $k$ is even also the apex-vertex in the centre. Hence every $k$-separator of $G$ is crossed by another $k$-separator. However, $G$ is far from $(k+1)$-connected, and in no way ‘basic’: every $42k$-clique can be replaced with an arbitrarily large $(k+1)$-connected graph, so $G$ represents a whole zoo of examples.
@@
Two tetra-separations $(A,B)$ and $(C,D)$ are nested if, after possibly exchanging the name $A$ with $B$ or the name $C$ with $D$, we have $A\subseteq C$ and $B\supseteq D$. A tetra-separation of $G$ is totally-nested if it is nested with every tetra-separation of $G$. Figure˜1.4 shows the totally-nested tetra-separations of an example-graph similar to the one shown in Figure˜1.3.
@@
In order to state our decomposition-result for 4-connected graphs, we need the following definitions to describe the ‘basic’ parts of the decomposition – those that are not quasi-5-connected. Graph-decompositions generalise tree-decompositions in that the decomposition-tree may take the form of an arbitrary graph, such as a cycle, in which case we speak of a cycle-decomposition [DJKK22]. Let $X$ be a graph. We call $X$ a generalised double-wheel with centre $\{u,v\}$ if $X$ is 4-connected and $u,v$ are distinct vertices of $X$ such that XuvX-u-v has a cycle-decomposition of adhesion 1 into $K_{2}$’s and triangles (Figure˜1.5). We call $X$ a cycle of $X_{1}$-torsos, …, $X_{k}$-torsos, $Y_{1}$-bags, …, $Y_{\ell}$-bags where the $X_{i}$ and $Y_{j}$ are graphs, if $X$ is 4-connected and has a cycle-decomposition with all adhesion-sets of size 2 such that for every bag either its torso is isomorphic to some $X_{i}$ or the bag itself is isomorphic to some $Y_{j}$ (Figure˜1.6). Recall that the only 3-connected graphs on $\leqslant 5$ vertices are $K_{4}$, the $4$-wheel, $K_{5}$ minus an edge, and $K_{5}$. A thickened $K_{4,m}$ is obtained from a $K_{4,m}$ by turning the left side into a $K_{4}$. A sprinkled $K_{4,m}$ is obtained from a $K_{4,m}$ by adding any number of edges to the left side.
@@
For 3-connected graphs, the strict tri-separation decomposition can be viewed as a refinement of the tri-separation decomposition, but the details are a bit technical; see Section˜15.
Algorithmic aspects. It is straightforward to come up with a polynomial-time algorithm that produces the decompositions in Theorem˜1 and ˜2 for all 4- and 3-connected graphs.
Korhonen [K24] recently showed that there is an algorithm that, given an $n$-vertex $m$-edge graph $G$ and an integer $k\geqslant 1$, returns a $k$-lean tree-decomposition of $G$ in time $k^{\mathcal{O}(k^{2})}n+\mathcal{O}(m)$. We believe that every 5-lean tree-decomposition of every 4-connected graph $G$ can be made into a tetra-separation decomposition of $G$ as in Theorem˜1 in linear time:
@@
There is an algorithm that, given a 4-connected $n$-vertex $m$-edge graph $G$, returns the tetra-separation decomposition of $G$ from Theorem˜1 in linear time $\mathcal{O}(n+m)$.
A positive answer would also provide a linear-time algorithm for the decomposition of 3-connected graphs in ˜2 by following the short proof of ˜2.
@@
Let $H$ be an arbitrary graph. The following definition of $H^{\Delta}$ is supported by Figure˜1.8. Let $U$ be the set of all vertices in $H$ of degree three. We obtain $H^{s}$ from $H$ by subdividing every edge that joins two vertices in $U$. Then we obtain $H^{s\Delta}$ from $H^{s}$ by making each neighbourhood $N_{H^{s}}(u)$ around a vertex $u\in U$ into a triangle $\Delta(u)$. Finally, we let $H^{\Delta}:=H^{s\Delta}-U$. If $H$ is quasi-4-connected and has >6>6 vertices, then $H^{\Delta}$ is 4-connected by Lemma˜12.2. The operation $H\mapsto H^{\Delta}$ is canonical: every isomorphism $H_{1}\to H_{2}$ induces an isomorphism $H^{\Delta}_{1}\to H^{\Delta}_{2}$.
@@
The operation $\tau\mapsto\tau^{\Delta}$ harmonises well with the other decomposition steps. Indeed, we have followed all steps to prove Figure˜1.8 below.
Vertex-transitive graphs. Recall that a graph $G$ is arc-transitive if for every two ordered pairs $(u_{i},v_{i})$ with $u_{i}v_{i}\in E(G)$ for $i=1,2$ there is an automorphism $\varphi$ of $G$ such that $\varphi(u_{1})=u_{2}$ and $\varphi(v_{1})=v_{2}$. Let $H$ be an $r$-vertex graph and let $G$ be an $r$-regular graph. An $H$-expansion of $G$ is a graph obtained from $G$ by replacing the vertices $v$ of $G$ with pairwise disjoint copies $H_{v}$ of $H$ and by replacing every edge $e=uv$ of $G$ with an $H_{u}$$H_{v}$ edge $\hat{e}$ so that $\hat{e}_{1}$ and $\hat{e}_{2}$ share no ends for every two $e_{1}\neq e_{2}\in E(G)$. Thus, the edges $\hat{e}$ form a perfect matching of the $H$-expansion. It was mostly222[CK23, Corollary 2] is slightly weaker than the result we mention here. The slight strengthening happens in Proposition 13.51 in this paper. shown in [CK23] that a finite connected graph $G$ is vertex-transitive if and only if $G$ is either
  • $\cdot$
    quasi-4-connected and vertex-transitive,
  • $\cdot$
    the $K_{3}$-expansion of a quasi-4-connected 4-regular arc-transitive graph,
  • $\cdot$
    a cycle, $K_{2}$ or $K_{1}$.
@@
A graph $G$ is 2-quasi-5-connected if $G$ is quasi-4-connected and every 4-separation $(A,B)$ of $G$ satisfies $|A\smallsetminus B|\leqslant 2$ or $|B\smallsetminus A|\leqslant 2$. For other definitions related to Figure˜1.8, see Section˜13.
{restatable}
mainresultmaintransitive A quasi-4-connected finite graph $G$ is vertex-transitive if and only if $G$ either is
  1. (i)
    a $(\geqslant 4)$-regular quasi-5-connected vertex-transitive graph,
  2. (ii)
    a 3-regular 2-quasi-5-connected vertex-transitive graph,
  3. (iii)
    a $K_{4}$-expansion of a quasi-5-connected 4-regular arc-transitive graph, or a vertex-transitive $C_{4}$-expansion of a quasi-5-connected 4-regular arc-transitive graph,
  4. (iv)
    a cycle of $K_{4}$-bags of length $\geqslant 4$ (Figure˜1.9),
  5. (v)
    a cycle alternating between $K_{4}$-bags and $C_{4}$-bags of length $\geqslant 4$ (Figure˜1.10),
  6. (vi)
    a cycle of triangle-bags of length $\geqslant 6$ (Figure˜1.12),
  7. (vii)
    a cycle alternating between $K_{2,2}$-bags and $C_{4}$-torsos of length $\geqslant 6$ (Figure˜1.11),
  8. (viii)
    a cycle of $C_{4}$-bags of length $\geqslant 4$ (Figure˜1.13) or $K_{3}\square K_{2}$,
  9. (ix)
    a cycle of $K_{2,2}$-bags of length $\geqslant 3$, the $K_{4}$-expansion thereof, or a vertex-transitive $C_{4}$-expansion thereof (Figure˜1.14), or
  10. (x)
    the line-graph of the cube, its $K_{4}$-expansion, or one of the three $C_{4}$-expansions of it depicted in Figure˜1.15.
Figure 1.9.
A cycle of $K_{4}$-bags
Figure 1.10.
A cycle alternating between $K_{4}$-bags and $C_{4}$-bags
Figure 1.11.
A cycle alternating between $K_{2,2}$-bags and $C_{4}$-torsos
Figure 1.12.
Cycles of triangle-bags
Figure 1.13.
Cycles of $C_{4}$-bags
Figure 1.14. A cycle of $K_{2,2}$-bags, its $K_{4}$-expansion, and two vertex-transitive $C_{4}$-expansions
Figure 1.15. The line-graph of the cube, its $K_{4}$-expansion, and its vertex-transitive $C_{4}$-expansions
@@ 6. $\cdot$
    Finite groups. The proof of Figure˜1.8 crucially relies on canonicity and total-nestedness. In a different direction, Stallings’ theorem applies tree-decomposition methods to Cayley graphs to detect product structure in groups. However, Stallings’ approach only works for infinite groups. Local separators have recently been used to prove a low-order Stallings-type theorem for finite nilpotent groups [CKKT24]. The proof heavily relies on the Tutte-type decomposition for local 2-separators [C22] and exploits the combination of canonicity and total-nestedness. The next step towards fully generalising Stallings theorem to finite groups will be to use versions for local separators of Theorem˜1 and of either ˜2 or [CK23].
@@ 1.1. Overview of the proof
In order to show that $\tau$ satisfies one of the outcomes of Theorem˜1, we will study how $\tau$ interacts with the tetra-separations of $G$. Indeed, the outcome will be determined by this. Let $(A,B)$ be a tetra-separation of $G$, which may possibly be crossed. We say that $(A,B)$ interlaces $\tau$ if, roughly speaking, both $A\smallsetminus B$ and $B\smallsetminus A$ intersect $\tau$. Here may think of $(A,B)$ as ‘cutting through’ $\tau$. Whether $\tau$ is interlaced or not already yields a first outcome:
@@
Assume now that $|Z|=0$. As our first step, we will use $(A,B)$ and $(C,D)$ to cut $G$ into four 2-connected pieces, arranged in the shape of a 4-cycle, and apply the Tutte-decomposition to each piece; see Figure˜1.16. Since $G$ is 4-connected, the Tutte-decompositions will be path-decompositions. We will then merge the four Tutte-decompositions into a cycle-decomposition $\mathcal{O}=(O,(G_{o})_{o\in V(O)})$. So all adhesion-sets of $\mathcal{O}$ have size two. We then use the 4-connectivity of $G$ to note that the torsos of $\mathcal{O}$ are triangles, 4-cycles, or 3-connected. Moreover, we will find some rules on how these types of torsos may be arranged on $O$. For instance, a torso that is a 4-cycle must always have two 3-connected torsos as neighbours, and if two triangle-torsos are neighbours then their adhesion-set is spanned by an edge.
@@
A major difference to tri-separations is that the connectivity-properties required of the separators are much more complex, as tetra-separations can cross in significantly more complicated ways than tri-separations; see Figure˜1.17 for an example.
@@
With ˜4 at hand, the problem of proving the hard backward implication of ˜3 becomes a problem about finding many paths in many configurations, roughly speaking. To this end, we develop a number of lemmata for finding such paths systematically, and then combine them to prove ˜3. Finding the paths takes a lot of work, especially when $O$ is short; see Figure˜1.18 for example.
@@
Proving ˜5, however, is a little tricky. We will assume for a contradicting that ˜5 fails, so some totally-nested tetra-separation $(E,F)$ will ‘cut through’ $\mathcal{O}$. Since $(E,F)$ is nested with both $(A,B)$ and $(C,D)$, we have $E\subseteq A\cap C$, say; see Figure˜1.19. We then start adjusting both $(A,B)$ and $(C,D)$: by moving bags of $\mathcal{O}$ from $A\smallsetminus B$ to $B\smallsetminus A$ and from $C\smallsetminus D$ to $D\smallsetminus C$, one bag at a time, while using the total-nestedness of $(E,F)$ to maintain the inclusion $E\subseteq A\cap C$. Eventually, there will be essentially only one bag $G_{o}$ left, and we will see that its torso $H_{o}$ is good. We will then show that $(E,F)=(U_{o},W_{o})$ is the only possibility. Since $(U_{o},W_{o})$ does not ‘cut through’ $\mathcal{O}$, this will yield a contradiction.

### v3

@@ Abstract.
As an application of the decomposition for 4-connectivity, we obtain a new theorem [KP26] characterising all vertex-transitive finite connected graphs as essentially quasi-5-connected or on a short explicit list of graphs.
@@
4-connected, decomposition, Tutte, canonical, double wheel, quasi-5-connected, 4-separation, 4-separator, tetra-separation, tetra-separator
@@ 1. Introduction
Background. To solve ˜1.1 for $k=0$ and $k=1$, one simply cuts along all $k$-separators of the graph. But already for $k=2$, this approach fails. Indeed, different 2-separators can cross in that they cut through each other, as happens in a cycle. Then if we cut at one of the 2-separators, we lose the other. Tutte realised that the solution is to cut the graph at all 2-separators that are totally-nested in that they are not crossed by any 2-separators, and indeed no conflicts arise when cutting in this way. The resulting pieces of Tutte’s decomposition are 3-connected, cycles or $K_{2}$’s. All decompositions for $k\leqslant 2$ can be described in terms of tree-decompositions. For instance, Tutte’s selection of 2-separators gives a recipe for a tree-decomposition with all adhesion-sets of size $2$ and all whose torsos [^1] A torso of a tree-decomposition is obtained from a bag by making each of its adhesion-sets into a clique. are 3-connected, cycles or $K_{2}$’s.
@@
There are $k$-connected graphs which just slightly fail to be $(k+1)$-connected, but all whose $k$-separators are crossed. For example, if $G$ is $k$-connected and the neighbourhoods of its vertices are precisely the $k$-separators of $G$, then the $k$-separators around neighbouring vertices cross, but $G$ fails to be $(k+1)$-connected. Such $G$ exist for all $k\geqslant 3$: take a circular saw as in fig.˜1.1, for instance. We believe that such graphs should not have to be decomposed any further at level $k$.
Figure $1.$ A circular saw has vertex-set $\mathbb{Z}_{n}\times\mathbb{Z}_{2}$ with $n\gg k$ and every vertex $(v,0)$ sends edges to the $k$ vertices $(v,1),\ldots,(v+k-1,1)$. It is $k$-connected, and when $k\geqslant 3$ its neighbourhoods of vertices are precisely its $k$-separators (lemma˜13.2).
@@
We consider the $k$-connected graph $G$ depicted in fig.˜1.2 if $k$ is odd and in fig.˜1.3 if $k$ is even. Every $k$-separator of $G$ is formed from a red set, a non-adjacent blue set, and if $k$ is even also the apex-vertex in the centre. Hence every $k$-separator of $G$ is crossed by another $k$-separator. However, $G$ is far from $(k+1)$-connected, and in no way ‘basic’: every $42k$-clique can be replaced with an arbitrarily large $(k+1)$-connected graph, so $G$ represents a whole zoo of examples.
@@
Two tetra-separations $(A,B)$ and $(C,D)$ are nested if, after possibly exchanging the name $A$ with $B$ or the name $C$ with $D$, we have $A\subseteq C$ and $B\supseteq D$. A tetra-separation of $G$ is totally-nested if it is nested with every tetra-separation of $G$. fig.˜1.4 shows the totally-nested tetra-separations of an example-graph similar to the one shown in fig.˜1.3.
@@
In order to state our decomposition-result for 4-connected graphs, we need the following definitions to describe the ‘basic’ parts of the decomposition – those that are not quasi-5-connected. Graph-decompositions generalise tree-decompositions in that the decomposition-tree may take the form of an arbitrary graph, such as a cycle, in which case we speak of a cycle-decomposition [RRPJ26]. Let $X$ be a graph. We call $X$ a generalised double-wheel with centre $\{u,v\}$ if $X$ is 4-connected and $u,v$ are distinct vertices of $X$ such that XuvX-u-v has a cycle-decomposition of adhesion 1 into $K_{2}$’s and triangles (fig.˜1.5). We call $X$ a cycle of $X_{1}$-torsos, …, $X_{k}$-torsos, $Y_{1}$-bags, …, $Y_{\ell}$-bags where the $X_{i}$ and $Y_{j}$ are graphs, if $X$ is 4-connected and has a cycle-decomposition with all adhesion-sets of size 2 such that for every bag either its torso is isomorphic to some $X_{i}$ or the bag itself is isomorphic to some $Y_{j}$ (fig.˜1.6). Recall that the only 3-connected graphs on $\leqslant 5$ vertices are $K_{4}$, the $4$-wheel, $K_{5}$ minus an edge, and $K_{5}$. A thickened $K_{4,m}$ is obtained from a $K_{4,m}$ by turning the left side into a $K_{4}$. A sprinkled $K_{4,m}$ is obtained from a $K_{4,m}$ by adding any number of edges to the left side.
@@
For 3-connected graphs, the strict tri-separation decomposition can be viewed as a refinement of the tri-separation decomposition, but the details are a bit technical; see section˜14.
Algorithmic aspects. It is straightforward to come up with a polynomial-time algorithm that produces the decompositions in Theorem˜1 and ˜2 for all 4- and 3-connected finite graphs.
Korhonen [K25] recently showed that there is an algorithm that, given an $n$-vertex $m$-edge finite graph $G$ and an integer $k\geqslant 1$, returns a $k$-lean tree-decomposition of $G$ in time $k^{\mathcal{O}(k^{2})}n+\mathcal{O}(m)$. We believe that every 5-lean tree-decomposition of every 4-connected finite graph $G$ can be made into a tetra-separation decomposition of $G$ as in Theorem˜1 in linear time:
@@
There is an algorithm that, given a 4-connected $n$-vertex $m$-edge finite graph $G$, returns the tetra-separation decomposition of $G$ from Theorem˜1 in linear time $\mathcal{O}(n+m)$.
A positive answer would also provide a linear-time algorithm for the decomposition of 3-connected finite graphs in ˜2 by following the short proof of ˜2.
@@
Let $H$ be an arbitrary graph. The following definition of $H^{\Delta}$ is supported by fig.˜1.8. Let $U$ be the set of all vertices in $H$ of degree three. We obtain $H^{s}$ from $H$ by subdividing every edge that joins two vertices in $U$. Then we obtain $H^{s\Delta}$ from $H^{s}$ by making each neighbourhood $N_{H^{s}}(u)$ around a vertex $u\in U$ into a triangle $\Delta(u)$. Finally, we let $H^{\Delta}:=H^{s\Delta}-U$. If $H$ is quasi-4-connected and has >6>6 vertices, then $H^{\Delta}$ is 4-connected by lemma˜12.2. The operation $H\mapsto H^{\Delta}$ is canonical: every isomorphism $H_{1}\to H_{2}$ induces an isomorphism $H^{\Delta}_{1}\to H^{\Delta}_{2}$.
@@
The operation $\tau\mapsto\tau^{\Delta}$ harmonises well with the other decomposition steps. Indeed, we have followed all steps to classify all vertex-transitive finite graphs of low connectivity [KP26].
  1. $\cdot$
@@ 7. $\cdot$
    Finite groups. Stallings’ theorem applies tree-decomposition methods to Cayley graphs to detect product structure in groups. However, Stallings’ approach only works for infinite groups. Local separators have recently been used to prove a low-order Stallings-type theorem for finite nilpotent groups [CKKT24]. The proof heavily relies on the Tutte-type decomposition for local 2-separators [C22] and exploits the combination of canonicity and total-nestedness. The next step towards fully generalising Stallings theorem to finite groups will be to use versions for local separators of Theorem˜1 and of either ˜2 or [CK23].
@@ 1.1. Overview of the proof
In order to show that $\tau$ satisfies one of the outcomes of Theorem˜1, we will study how $\tau$ interacts with the tetra-separations of $G$. Indeed, the outcome will be determined by this. Let $(A,B)$ be a tetra-separation of $G$, which may possibly be crossed. We say that $(A,B)$ interlaces $\tau$ if, roughly speaking, both $A\smallsetminus B$ and $B\smallsetminus A$ intersect $\tau$. Here we may think of $(A,B)$ as ‘cutting through’ $\tau$. Whether $\tau$ is interlaced or not already yields a first outcome:
@@
Assume now that $|Z|=0$. As our first step, we will use $(A,B)$ and $(C,D)$ to cut $G$ into four 2-connected pieces, arranged in the shape of a 4-cycle, and apply the Tutte-decomposition to each piece; see fig.˜1.9. Since $G$ is 4-connected, the Tutte-decompositions will be path-decompositions. We will then merge the four Tutte-decompositions into a cycle-decomposition $\mathcal{O}=(O,(G_{o})_{o\in V(O)})$. So all adhesion-sets of $\mathcal{O}$ have size two. We then use the 4-connectivity of $G$ to note that the torsos of $\mathcal{O}$ are triangles, 4-cycles, or 3-connected. Moreover, we will find some rules on how these types of torsos may be arranged on $O$. For instance, a torso that is a 4-cycle must always have two 3-connected torsos as neighbours, and if two triangle-torsos are neighbours then their adhesion-set is spanned by an edge.
@@
A major difference to tri-separations is that the connectivity-properties required of the separators are much more complex, as tetra-separations can cross in significantly more complicated ways than tri-separations; see fig.˜1.10 for an example.
@@
With ˜4 at hand, the problem of proving the hard backward implication of ˜3 becomes a problem about finding many paths in many configurations, roughly speaking. To this end, we develop a number of lemmata for finding such paths systematically, and then combine them to prove ˜3. Finding the paths takes a lot of work, especially when $O$ is short; see fig.˜1.11 for example.
@@
Proving ˜5, however, is a little tricky. We will assume for a contradiction that ˜5 fails, so some totally-nested tetra-separation $(E,F)$ will ‘cut through’ $\mathcal{O}$. Since $(E,F)$ is nested with both $(A,B)$ and $(C,D)$, we have $E\subseteq A\cap C$, say; see fig.˜1.12. We then start adjusting both $(A,B)$ and $(C,D)$: by moving bags of $\mathcal{O}$ from $A\smallsetminus B$ to $B\smallsetminus A$ and from $C\smallsetminus D$ to $D\smallsetminus C$, one bag at a time, while using the total-nestedness of $(E,F)$ to maintain the inclusion $E\subseteq A\cap C$. Eventually, there will be essentially only one bag $G_{o}$ left, and we will see that its torso $H_{o}$ is good. We will then show that $(E,F)=(U_{o},W_{o})$ is the only possibility. Since $(U_{o},W_{o})$ does not ‘cut through’ $\mathcal{O}$, this will yield a contradiction.
@@
Mixed-separators (without additional conditions on their vertices) have been used before, see for example Mader’s theorem [M78, R24] or [BH67, JKS24, HHSS20].